OpenStudy (anonymous):

(3x^3y)^2(-4x^2y^4)^3

OpenStudy (anonymous):

(please rewrite with parentheses where needed. such as( 3x^3)*y)^2

OpenStudy (zehanz):

Just pasting the formula into the equation editor makes it look better already:\[(3x^3y)^2(-4x^2y^4)^3 \]Use the following three rules: \[(ab)^c=a^cb^c\]\[(a^b)^c=a^{bc}\]\[a^ba^c=a^{b+c}\]You'll need all three of them, in this order! First rule: distribute the exponents outside of the brackets to each of the factors within. Second rule: power-of-power means: multiply the exponents. Third rule: look for powers with same base, write as one with sum of exponents. Done!